1.The rectangular channel has a width of 4 m and transports water at 1.20(106) liters/min. If the elevation of the bed drops by 0.4 ft, determine the depth y2 after the depression.


2. A flow of 60 m3/s passes under the sluice gate. If y1 = 6 in. determine the depth y2 and the type of flow downstream. Also, what is the maximum volumetric flow that can pass through the gate?


1
Expert's answer
2022-01-18T15:13:06-0500

1.

"y_2=y_1+\frac{\dot V l^2}2=6.7~m,"

2.

"y_2=y_1-\frac{\dot V l^2}2=4~m,"

"\dot V_{max}=\dot V\frac{y_1+y_2}{y_1}=100~\frac{m^3}s."